Suppose you are working in the financial management team at Madison Square Garden and need to compare two capital projects to identify a better one.

●Project 1 (upgrading seats) has an expected useful life of 10 years and anticipated annual cash flows of $90,000. The initial cost is $520,000.

●Project 2 (replacing scoreboard) has an expected useful life of 7 years and anticipated annual cash flows of $80,000. The initial cost is $250,000.

●Your organization will use bank loan with 10% annual interest rate.

Q1. Use methods of NPV and IRR and identify which of projects should be accepted

Q2. What would happen if annual interest rate is 13% and 15%?
Answer #1

C F G H J K L M N O Year WN PV Factor@10% Project 1 PV of Project 1 Project 2 PV of Project 2 11 $-5,20,000.00 $-5,20,000.00Year PV Factor@15% Project 1 PV of Project 1 Project 2 PV of Project 2 11 $-5,20,000.00 $-5,20,000.00 $-2,50,000.00 $ -2,50,0Year OMN00 PV Factor@10% =1/1.1^B3 =1/1.1^B4 =1/1.1^B5 =1/1.1^B6 =1/1.1^B7 =1/1.1^B8 =1/1.1^B9 =1/1.1^B10 =1/1.1^B11 =1/1.1^B